OpenStudy (anonymous):

Write an equation of the line perpendicular to y = -4x – 3 that contains the point (7, 3).

OpenStudy (anonymous):

Hint: if the lines are Perpendicular their Slopes m1*m2 = -1

OpenStudy (ajprincess):

The gradient of the given line=\(m_1\) The gradient of the required line=\(m\) \(m*m_1=-1\) \(m_1=-4\) \(m*-4=-1\) \(m=\large\frac{-1}{-4}\) \(m=\large\frac{1}{4}\) \(m=\large\frac{y-y_1}{x-x_1}\) \(\large\frac{1}{4}=\frac{y-3}{x-7}\) Can u do nw?

OpenStudy (anonymous):

@Yahoo! Haha thanx Yahoo! again. lol. @ajprincess yea. thank you!

OpenStudy (anonymous):

Welcome...)

OpenStudy (ajprincess):

yw:)